In a meeting, Matt estimates the average miles per gallon (mpg) of the fleet cars is 24 mpg. With a certain level of confidence, he announces the actual mpg is in the 21 to 27 mpg range. What is the Margin of Error?

A. 6 mpg

B. 24 mpg

C. 3 mpg

D. 0.50 mpg
Answer #1

Solution :

  = (Lower confidence interval + Upper confidence interval ) / 2

Sample mean = = (21+27) / 2=24

Margin of error = E = Upper confidence interval -   = 27-24=3

Margin of error = E=3mpg